OpenStudy (anonymous):

lee withdrew $200 from his account and asked for $5 and $10 bills. He received two more $10 bills than $5 bills. how many bills did he receive in all.

OpenStudy (aripotta):

system of equations: 5a + 10b = 200 b = a + 2 substitute: 5a + 10(a + 2) = 200 distribute: 5a + 10a + 20 = 200 combine like terms: 15a + 20 = 200 subtract 20 from both sides: 15a = 180 divide by 15 on both sides: a = 12 substitute to solve for b: b = 12 + 2 b = 14 add them together: 14 + 12 = 26 so he received 26 bills

OpenStudy (aripotta):

does that make sense?

OpenStudy (anonymous):

I got the same as AriPotta, so his should be right, and it makes sense to me.

OpenStudy (aripotta):

@gkendryna do you understand what i did?